OpenStudy (anonymous):

What is the solution to the equation fraction 4 over 5 n minus fraction 3 over 5 equals fraction 1 over 5 n

OpenStudy (anonymous):

n = 1

OpenStudy (harley1312):

N=1

OpenStudy (ddcamp):

Is this the equation? \[\frac{ 4 }{ 5n } - \frac{ 3 }{ 5 } = \frac{ 1 }{ 5n }\]

OpenStudy (ddcamp):

Let's start by isolating the variable: